OpenStudy (anonymous):

For the following system, if you isolated y in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 2 x − y = −6 −6 − x −6 + x 6 − x 6 + x

OpenStudy (anonymous):

@yummydum

OpenStudy (anonymous):

solve \(x − y = −6\) for \(y\)

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

you're not trying...

OpenStudy (anonymous):

solve for y...its basic algebra bro, c'mon, i know you know this

OpenStudy (anonymous):

oh! I was trying! I understood differently.

OpenStudy (anonymous):

here's an example of how i would solve for x:\[x-y=-6\]add y to both sides:\[x=y-6\] so just solve for y:\[x-y=-6\]then...?

OpenStudy (anonymous):

okay... I got this..

OpenStudy (anonymous):

just get it to be y=... i

OpenStudy (anonymous):

I'm still not sure, I'm thinking C.

OpenStudy (anonymous):

close, it's D

OpenStudy (anonymous):

damn! I thought I had it lol. well thanks anyways.

OpenStudy (anonymous):

I definitely am...

OpenStudy (anonymous):

whatever bro, i tried explaining the deal to you but you're not wanting to understand....you're on your own, man..

OpenStudy (anonymous):

wow, thank God you're not a teacher.